# Finely Stratified Rerandomization Designs

### v1

@@ Abstract
We study estimation and inference on causal parameters under finely stratified rerandomization designs, which use baseline covariates to match units into groups (e.g. matched pairs), then rerandomize within-group treatment assignments until a balance criterion is satisfied. We show that finely stratified rerandomization does partially linear regression adjustment “by design,” providing nonparametric control over the covariates used for stratification, and linear control over the rerandomization covariates. We also introduce novel rerandomization criteria, allowing for nonlinear imbalance metrics and proposing a minimax scheme that optimizes the balance criterion using pilot data or prior information provided by the researcher. While the asymptotic distribution of generalized method of moments (GMM) estimators under stratified rerandomization is generically non-Gaussian, we show how to restore asymptotic normality using optimal ex-post linear adjustment. This allows us to provide simple asymptotically exact inference methods for superpopulation parameters, as well as efficient conservative inference methods for finite population parameters.
@@ 1 Introduction
Stratified randomization is commonly used to increase statistical precision in experimental research. [^2] For example, Cytrynbaum, (2023) reports a survey of 50 experimental papers in the AER and AEJ from 2018-2023, where 57% used some form of stratified randomization. Recent theoretical work (e.g. Bai et al., (2021)) has shown that fine stratification, which randomizes within small groups of units tightly matched on baseline covariate information, makes unadjusted estimators like difference of means semiparametrically efficient. [^3] See Cytrynbaum, (2024), Armstrong, (2022), and Bai et al., (2024) for more detailed discussion. In finite samples, however, the performance of such designs can deteriorate rapidly with the dimension of the stratification variables due to a curse of dimensionality in matching. [^4] Under regularity conditions, the convergence rate of finite sample variance to asymptotic variance is $O(n^{-2/(d+1)})$ for dimension $d$ covariates, see Cytrynbaum, (2024). This motivates the search for alternative designs that insist upon nonparametric balance for a few important covariates, but only attempt to balance linear functions of the remaining variables. In this paper, we propose finely stratified rerandomization designs, which first tightly match the units into groups using a small set of important covariates, then rerandomize within groups until a balance criterion on the remaining covariates is satisfied.
@@
Our first contribution is to derive the asymptotic distribution of generalized method of moments (GMM) estimators under stratified rerandomization, allowing estimation and inference on generic causal parameters defined by moment equalities. We consider both superpopulation and finite population parameters, the latter of which may be more appropriate for experiments run in a convenience sample (Abadie et al., (2014)). As in previous work on rerandomization (e.g. Li et al., (2018)), the asymptotic distribution of GMM estimators is an independent sum of a normal and a truncated normal term. Modulo this residual truncated term, we show that the asymptotic variance of unadjusted estimation under stratified rerandomization is the same as that of semiparametrically adjusted GMM (e.g. Graham, (2011)) under an iid design. Intuitively, stratified rerandomization implements partially linear regression adjustment “by design.”
@@
Our second contribution is to introduce several novel forms of rerandomization based on nonlinear balance criteria. For example, we allow acceptance or rejection based on the difference of covariate density estimates within each treatment arm, attempting to balance nonlinear features of the covariate distribution. Similarly, we propose a design that rerandomizes until a nonlinear estimate of the propensity score is approximately constant, effectively forcing the covariates to have no predictive power for treatment assignments. In both cases, these nonlinear rerandomization schemes are asymptotically equivalent to standard rerandomization based on a difference of covariate means, but with an implicit choice of covariates and acceptance region, which we characterize.
@@
Our third contribution is to study optimization of the balance criterion itself. We suggest a novel minimax approach that allows the researcher to specify prior information about the relationship between covariates and outcomes, then rerandomizes until the worst case correlation consistent with this prior information is small. If the prior information set contains the truth, this design provides strong control over the variance of the truncated normal term in the asymptotic distribution. Building on this, we show that if the prior information set is a confidence region estimated from pilot data, then this minimax design bounds the truncated normal variance with high probability.
@@
Our fourth contribution is to provide simple t-statistic based inference methods for general causal parameters under stratified rerandomization designs. To do this, we first characterize and provide a feasible implementation of the optimal ex-post linear adjustment for GMM estimation under stratified rerandomization. [^5] This extends recent work on optimal adjustment under pure stratified randomization for ATE estimation, e.g. see Cytrynbaum, (2023), Bai et al., (2023), or Liu and Yang, (2020). Crucially, optimal ex-post adjustment makes the asymptotic distribution insensitive to the rerandomization acceptance criterion, removing the truncated normal term from the limiting distribution and restoring asymptotic normality. For superpopulation parameters, our inference methods are asymptotically exact. For finite population parameters, our methods are asymptotically conservative, but still exploit the efficiency gains from both stratified rerandomization and ex-post optimal adjustment.
@@ 1.1 Related Literature
For our main asymptotic theory (Section 3), the most closely related previous work is Wang et al., (2021) and Bai et al., (2024). Wang et al., (2021) study estimation of the sample average treatment effect (SATE) under stratified rerandomization, with quadratic imbalance metrics based on the Mahalanobis norm. We study rerandomization within data-adaptive fine strata, providing asymptotic theory for generic superpopulation and finite population causal parameters defined by moment equalities. We also allow for essentially arbitrary rerandomization acceptance criteria, not necessarily based on quadratic forms. Bai et al., (2024) study estimation of superpopulation parameters defined by moment equalities under pure stratified randomization. We extend these results to stratified rerandomization as well as generic finite population parameters, providing “SATE-like” versions of the parameters in Bai et al., (2024). [^6] These parameters can be seen as causal versions of the conditional estimand defined in Abadie et al., (2014). In concurrent work, Wang and Li, (2024) study GMM estimation of univariate superpopulation parameters under stratified rerandomization with fixed, discrete strata. We study significantly more general forms of stratification and rerandomization criteria than considered in their work, allowing for both finite and superpopulation parameters of arbitrary fixed dimension.
@@ 2 Framework and Designs
    (Stratification). Partition the experimental units into $n/k$ disjoint groups $g$ with $\{1,\dots,n\}=\bigcup_{g}g$ disjointly and $|g|=k$. Let $\psi=\psi(R)$ with $\psi\in\mathbb{R}^{d_{\psi}}$ denote a vector of stratification variables, which may be continuous or discrete. Suppose the groups satisfy the homogeneity condition [^7] The matching condition in Equation 2.1 was introduced by Bai et al., (2021) for matched pairs randomization ($k=2$). See Bai, (2022) and Cytrynbaum, (2024) for generalizations.
    $\frac{1}{n}\sum_{g}\sum_{i,j\in g}|\psi_{i}-\psi_{j}|_{2}^{2}=o_{p}(1).$ (2.1)
    Require that the groups only depend on the stratification variables $\psi_{1:n}$ and data-independent randomness $\pi_{n}$, so that $g=g(\psi_{1:n},\pi_{n})$ for each $g$.
@@
    (Randomization). Independently for each $|g|=k$, draw treatment variables $(D_{i})_{i\in g}$ by setting $D_{i}=1$ for exactly $l$ out of $k$ units, uniformly at random.
@@
Stratification without rerandomization can be obtained by setting $A=\mathbb{R}^{d_{h}}$ in Definition 2.1. Treatment effect estimation under such designs was studied in Bai, (2022), Cytrynbaum, (2024), and Bai et al., (2024). Definition 2.1 allows for fine stratification (also known as matched k-tuples), with the number of data-dependent groups $g=g(\psi_{1:n},\pi_{n})$ growing with $n$. It also allows for coarse stratification with fixed strata $s\in\{1,\dots,m\}$ and fixed $m$, as in Bugni et al., (2018), which can be obtained in this framework by setting $\psi=s$ and matching units into groups at random within the strata $\{i:s_{i}=k\}$.
@@
For $p=l/k$, we say that $D_{1:n}$ are completely randomized with probability $p$ if $P(D_{1:n}=d_{1:n})=1/\binom{n}{np}$ for all $d_{1:n}$ with $\sum_{i}d_{i}=np$ [^10] For notational simplicity, we may assume that $n=lk$ for some $l\in\mathbb{N}$. If so, we denote $D_{1:n}\sim\operatorname{CR}(p)$. Cytrynbaum, (2024) shows that $\operatorname{CR}(p)$ randomization can be obtained by setting $\psi=1$ and $A=\mathbb{R}^{d_{h}}$ in Definition 2.1, matching units into groups at random.
Causal Estimands. Next, we introduce a generic family of causal estimands defined by moment equalities. Let $g(D,R,S,\theta)\in\mathbb{R}^{d_{g}}$ be a score function for generalized method of moments (GMM) estimation. Recall $W=(R,S(1),S(0))$ and define the projection $\phi(W,\theta)=E[g(D,R,S,\theta)|W]$, so that $E[\phi(W,\theta)]=0$ $\iff$ $E[g(D,R,S,\theta)]=0$.
@@
This is the first order condition of a treatment effect prediction problem in the complier population. In particular, for $\tau\equiv Y(1)-Y(0)$, the parameter $\theta_{0}$ is the best parametric predictor of treatment effects for compliers:
@@ 3 Asymptotics for GMM Estimation
In this section, we characterize the asymptotic distribution of the GMM estimator $\widehat{\theta}$ under stratified rerandomization designs, as in Definition 2.1. We show that the variance under stratified rerandomization is proportional to the residuals of a partially linear regression model, up to a rerandomization imbalance term. In this sense, stratified rerandomization does partially linear regression adjustment “by design.” First, we state some technical regularity conditions that are needed for the following results.
@@
Suppose $A\in\mathbb{R}^{d_{h}}$ has non-empty interior and $\operatorname{Leb}(\partial A)=0$, [^13] Note that $\partial A$ denotes the boundary of $A$, the limit points of both $A$ and $A^{c}$. and require $E[\operatorname{Var}(h|\psi)]\succ 0$ and $E[|\psi|_{2}^{2}+|h|_{2}^{2}]<\infty$.
Next we state the technical conditions needed for GMM estimation. Define the matrix $G=E[(\partial/\partial\theta^{\prime})\phi(W,\theta)]|_{\theta=\theta_{0}}\in\mathbb{R}^{d_{g}\times d_{\theta}}$ and let $g_{d}(W,\theta)=g(d,R,S(d),\theta)$ for $d\in\{0,1\}$. Recall that $|B|_{F}^{2}=\sum_{ij}B_{ij}^{2}$ for any matrix $B$.
@@
    We have $E[g_{d}(W,\theta_{0})^{2}]<\infty$ and $E[\sup_{\theta\in\Theta}|g_{d}(W,\theta)|_{2}]<\infty$. Also $\theta\to g_{d}(W,\theta)$ is continuous almost surely, and $\Theta$ is compact with $\theta_{0}$ in the interior.
@@
    The matrix $G$ is full rank, and $g_{0}(\theta)=0$ iff $\theta=\theta_{0}$.
@@
Orthogonal Expansion. Recall $\phi(W,\theta)=E[g(D,R,S,\theta)|W]$ for $W=(R,S(1),S(0))$. Define the assignment influence component $\delta(W,\theta)\equiv\operatorname{Var}(D)(g_{1}(W,\theta)-g_{0}(W,\theta))$. For the Horvitz-Thompson weights $H=(D-p)/(p-p^{2})$, a simple calculation shows that we can expand
$g(D,R,S,\theta)=\phi(W,\theta)+H\delta(W,\theta).$ (3.1)
Our work below shows that $\delta(W,\theta)$ parameterizes estimator variance due to assignment, while $\phi(W,\theta)$ parameterizes variance due to random sampling. In what follows, we work directly with this expansion.
@@
Continuing Example 2.6 above, let $\bar{Y}=(1-p)Y(1)+pY(0)$, a convex combination that summarizes each unit’s potential outcome level. Then for the score $g(D,Y,\theta)=HY-\theta$, we have $\delta(W,\theta)=\bar{Y}$ and
$E_{n}[H_{i}\delta(W_{i})]=\bar{\delta}_{1}-\bar{\delta}_{0}=\frac{\operatorname{Cov}_{n}(D_{i},\bar{Y}_{i})}{\operatorname{Var}_{n}(D_{i})}.$
Intuitively, $E_{n}[H_{i}\delta(W_{i})]$ from Equation 3.1 isolates the estimator variance due to chance correlations between the assignments $D_{i}$ and outcome levels $\bar{Y}_{i}$. By contrast, $\phi(W,\theta)=Y(1)-Y(0)-\theta$ does not depend on assignments $D_{i}$, and $\operatorname{Var}(\phi(W,\theta))=\operatorname{Var}(Y(1)-Y(0))$ isolates the estimator variance due to random sampling of treatment effects.

### v2

@@ Abstract
We study estimation and inference on causal parameters under finely stratified rerandomization designs, which use baseline covariates to match units into groups (e.g. matched pairs), then rerandomize within-group treatment assignments until a balance criterion is satisfied. We show that finely stratified rerandomization does partially linear regression adjustment “by design,” providing nonparametric control over the stratified covariates and linear control over the rerandomized covariates. We introduce several new rerandomization designs, allowing for imbalance metrics based on nonlinear estimators. We also propose a novel minimax scheme that uses pilot data or prior information to minimize the computational cost of rerandomization, subject to a strict bound on statistical efficiency. While the asymptotic distribution of generalized method of moments (GMM) estimators under stratified rerandomization is generically non-normal, we show how to restore asymptotic normality using ex-post linear adjustment tailored to the stratification. This enables simple asymptotically exact inference on superpopulation parameters, as well as efficient conservative inference on finite population parameters.
@@ 1 Introduction
Stratified randomization is commonly used to increase statistical precision in experimental research. [^2] For example, Cytrynbaum, (2023) reports a survey of 50 experimental papers in the AER and AEJ from 2018-2023, where 57% used some form of stratified randomization. Recent theoretical work (e.g. Bai et al., (2021)) has shown that fine stratification, which randomizes within small groups of units tightly matched on baseline covariate information, makes unadjusted estimators like difference of means semiparametrically efficient. [^3] See Cytrynbaum, (2024), Armstrong, (2022), and Bai et al., (2024) for more detailed discussion. In finite samples, however, the performance of such designs can deteriorate rapidly with the dimension of the stratification variables due to a curse of dimensionality in matching. [^4] Under regularity conditions, the convergence rate of finite sample variance to asymptotic variance is $O(n^{-2/(d+1)})$ for dimension $d$ covariates, see Cytrynbaum, (2024). This motivates the search for alternative designs that insist upon nonparametric balance for a few important covariates, but only attempt to balance linear functions of the remaining variables. In this paper, we study finely stratified rerandomization designs, which first tightly match the units into groups using a small set of important covariates, then rerandomize within-groups treatment assignments until a balance criterion on the remaining covariates is satisfied.
@@
Our first contribution is to derive the asymptotic distribution of generalized method of moments (GMM) estimators under stratified rerandomization, allowing for estimation of generic causal parameters defined by moment equalities. We consider both superpopulation and finite population parameters, the latter of which may be more appropriate for experiments run in a convenience sample (Abadie et al., (2014)). As in previous work on rerandomization (e.g. Li et al., (2018)), the asymptotic distribution of GMM estimators is an independent sum of a normal and a truncated normal term. We show that, modulo this residual truncated term, the asymptotic variance of unadjusted estimation under stratified rerandomization is the same as that of semiparametrically adjusted GMM (e.g. Graham, (2011)) under an iid design. Intuitively, stratified rerandomization implements partially linear regression adjustment “by design.”
@@
Our second contribution is to introduce several novel forms of rerandomization based on nonlinear balance criteria. For example, we allow acceptance or rejection of an allocation based on the difference of covariate density estimates within each treatment arm, attempting to balance nonlinear features of the covariate distribution. Similarly, we propose a design that rerandomizes until a nonlinear estimate of the propensity score is approximately constant, effectively forcing the covariates to have no predictive power for treatment assignments. In both cases, these nonlinear rerandomization schemes are asymptotically equivalent to standard rerandomization based on a difference of covariate means, but with an implicit choice of covariates and acceptance region, which we characterize.
@@
Our third contribution is to study optimization of the balance criterion itself. We propose a novel minimax approach that allows the researcher to specify prior information about the relationship between covariates and outcomes, then rerandomizes until the worst case correlation between treatments and covariates consistent with this prior is small. We prove that this design minimizes the (asymptotic) computational cost of rerandomization, subject to a strict bound on statistical efficiency over the set of DGP’s consistent with the prior. If the prior information set contains the truth, this design strictly bounds the asymptotic variance within a small additive factor of the optimal semiparametrically adjusted variance. Extending this result, we show that if the prior information set is a confidence region estimated from pilot data, then this minimax design bounds the asymptotic variance in the main experiment with high probability.
@@
Our fourth contribution is to provide simple t-statistic and Wald based inference methods for general causal parameters under stratified rerandomization designs. To do this, we first characterize and provide a feasible implementation of the optimal ex-post linear adjustment for GMM estimation under stratified rerandomization. [^5] This extends recent work on optimal adjustment under pure stratified randomization for ATE estimation, e.g. see Cytrynbaum, (2023), Bai et al., (2023), or Liu and Yang, (2020). Crucially, optimal ex-post adjustment makes the asymptotic distribution insensitive to the rerandomization acceptance criterion, removing the truncated normal term from the limiting distribution and restoring asymptotic normality. For superpopulation parameters, our inference methods are asymptotically exact. For finite population parameters, our inference is conservative due to non-identification of the asymptotic variance, but still exploits the efficiency gains from both stratified rerandomization and ex-post optimal adjustment.
@@ 1.1 Related Literature
For our main asymptotic theory (Section 3), the most closely related previous work is Wang et al., (2021) and Bai et al., (2024). Wang et al., (2021) study estimation of the sample average treatment effect (SATE) under stratified rerandomization, with quadratic imbalance metrics based on the Mahalanobis norm. We study rerandomization within data-adaptive fine strata, providing asymptotic theory for generic superpopulation and finite population causal parameters defined by moment equalities. We also allow for essentially arbitrary rerandomization acceptance criteria, not necessarily based on quadratic forms. Bai et al., (2024) study estimation of superpopulation parameters defined by moment equalities under pure stratified randomization. We extend these results to stratified rerandomization as well as generic finite population parameters, providing “SATE-like” versions of the parameters in Bai et al., (2024). [^6] These parameters can be seen as causal versions of the conditional estimand defined in Abadie et al., (2014). In concurrent work, Wang and Li, (2024) study GMM estimation of univariate superpopulation parameters under stratified rerandomization with fixed, discrete strata. We study significantly more general forms of stratification and rerandomization criteria than considered in their work, allowing for both finite and superpopulation parameters of arbitrary dimension and fine stratification with continuous covariates.
@@ 2 Framework and Designs
    (Stratification). Partition the experimental units into $n/k$ disjoint groups $s$ with $\{1,\dots,n\}=\bigcup_{s}s$ disjointly and $|s|=k$. Let $\psi=\psi(R)$ with $\psi\in\mathbb{R}^{d_{\psi}}$ denote a vector of stratification variables, which may be continuous or discrete. Suppose the groups satisfy the homogeneity condition [^7] The matching condition in Equation 2.1 was introduced by Bai et al., (2021) for matched pairs randomization ($k=2$). See Bai, (2022) and Cytrynbaum, (2024) for generalizations.
    $\frac{1}{n}\sum_{s}\sum_{i,j\in s}|\psi_{i}-\psi_{j}|_{2}^{2}=o_{p}(1).$ (2.1)
    Require that the groups only depend on the stratification variables $\psi_{1:n}$ and data-independent randomness $\pi_{n}$, so that $s=s(\psi_{1:n},\pi_{n})$ for each $s$.
@@
    (Randomization). Independently for each $|s|=k$, draw treatment variables $(D_{i})_{i\in s}$ by setting $D_{i}=1$ for exactly $l$ out of $k$ units, uniformly at random.
@@
Stratification without rerandomization can be obtained by setting $A=\mathbb{R}^{d_{h}}$ in Definition 2.1. Treatment effect estimation under such designs was studied in Bai, (2022), Cytrynbaum, (2024), and Bai et al., (2024). Definition 2.1 allows for fine stratification (also known as matched k-tuples), with the number of data-dependent groups $s=s(\psi_{1:n},\pi_{n})$ growing with $n$. It also allows for coarse stratification with fixed strata $T\in\{1,\dots,m\}$ and fixed $m$, as in Bugni et al., (2018), which can be obtained in this framework by setting $\psi=T$ and matching units into groups $s$ at random within each $\{i:T_{i}=k\}$.
@@
For $p=l/k$, we say that $D_{1:n}$ are completely randomized with probability $p$ if $P(D_{1:n}=d_{1:n})=1/\binom{n}{np}$ for all $d_{1:n}$ with $\sum_{i}d_{i}=np$ [^10] For notational simplicity, we may assume that $n=lk$ for some $l\in\mathbb{N}$. If so, we denote $D_{1:n}\sim\operatorname{CR}(p)$. Cytrynbaum, (2024) shows that $\operatorname{CR}(p)$ randomization can be obtained by setting $\psi=1$ and $A=\mathbb{R}^{d_{h}}$ in Definition 2.1, matching units into groups at random. Intuitively, random matched k-tuples is equivalent to complete randomization.
Causal Estimands. Next, we introduce a generic family of causal estimands defined by moment equalities. Let $g(D,R,S,\theta)\in\mathbb{R}^{d_{g}}$ be a score function for generalized method of moments (GMM) estimation. Recall $W=(R,S(1),S(0))$ and for $D|W\sim\operatorname{Bernoulli}(p)$ define $\phi(W,\theta)=E[g(D,R,S,\theta)|W]=pg(1,R,S(1),\theta)+(1-p)g(0,R,S(0),\theta)$. By construction, we have $E[\phi(W,\theta)]=0$ $\iff$ $E[g(D,R,S,\theta)]=0$. The function $\phi(W,\theta)$ provides a convenient parameterization to introduce our causal estimands.
@@
Then $E[\phi(W,\theta)]=0$ is the first order condition of a treatment effect prediction problem in the complier population. In particular, for $\tau\equiv Y(1)-Y(0)$, the parameter $\theta_{0}$ is the best parametric predictor of treatment effects for compliers:
@@
Remark 2.9 (M-estimation).
Our results below also extend to M-estimators of the form $\widehat{\theta}=\operatorname*{argmax}_{\theta}E_{n}[m(D_{i},R_{i},S_{i},\theta)]$, even when $\widehat{\theta}$ cannot be formulated as a GMM estimator e.g. due to the existence of local maxima. For example, this happens in some nonconvex problems in density estimation (Newey and McFadden, (1994)). We briefly discuss this extension in Section 8.4.
@@ 3 Asymptotics for GMM Estimation
In this section, we characterize the asymptotic distribution of the GMM estimator $\widehat{\theta}$ under stratified rerandomization designs, as in Definition 2.1. We show that the variance under stratified rerandomization is proportional to the residuals of a partially linear regression model, up to an extra term that reflects slackness in the rerandomization criterion. In this sense, stratified rerandomization does partially linear regression adjustment “by design.” First, we state some technical conditions that are needed for the following results.
@@
Suppose $A\subseteq\mathbb{R}^{d_{h}}$ has non-empty interior and $\operatorname{Leb}(\partial A)=0$, [^13] Note that $\partial A$ denotes the boundary of $A$, the limit points of both $A$ and $A^{c}$. and require $E[\operatorname{Var}(h|\psi)]\succ 0$ and $E[|\psi|_{2}^{2}+|h|_{2}^{2}]<\infty$.
Next we state the technical conditions needed for GMM estimation. Define the matrix $G=E[(\partial/\partial\theta^{\prime})\phi(W,\theta)]|_{\theta=\theta_{0}}\in\mathbb{R}^{d_{g}\times d_{\theta}}$ and let $g_{d}(W,\theta)=g(d,R,S(d),\theta)$ for $d\in\{0,1\}$. Recall the Frobenius norm $|B|_{F}^{2}=\sum_{ij}B_{ij}^{2}$ for any matrix $B$.
@@
    (Identification). The matrix $G$ is full rank, and $g_{0}(\theta)=0$ iff $\theta=\theta_{0}$.
@@
    We have $E[g_{d}(W,\theta_{0})^{2}]<\infty$ and $E[\sup_{\theta\in\Theta}|g_{d}(W,\theta)|_{2}]<\infty$. Also $\theta\to g_{d}(W,\theta)$ is continuous almost surely, and $\Theta$ is compact. [^14] We can formally resolve measurability issues with the sup expressions by either (1) explicitly working with outer probability (e.g. van der Vaart and Wellner, (1996)) or (2) requiring that $\{g_{d}(\cdot,\theta),\theta\in\Theta\}$ is universally separable for $d=0,1$ (Pollard, (1984), p.38). To focus on the practical design issues, we avoid this formalism, implicitly assuming that all quantities are appropriately measurable.
@@
Orthogonal Expansion. Recall $\phi(W,\theta)=E[g(D,R,S,\theta)|W]$ for $W=(R,S(1),S(0))$. Define the assignment influence component $a(W,\theta)\equiv\operatorname{Var}(D)(g_{1}(W,\theta)-g_{0}(W,\theta))$. For Horvitz-Thompson weights $H=(D-p)/(p-p^{2})$, a simple calculation shows that we can expand
$g(D,R,S,\theta)=\phi(W,\theta)+Ha(W,\theta).$ (3.1)
Our work below shows that $a(W,\theta)$ parameterizes estimator variance due to random assignment, while $\phi(W,\theta)$ parameterizes the variance due to random sampling. We work directly with this expansion in what follows.
@@
Continuing Example 2.6 above, let $\bar{Y}=(1-p)Y(1)+pY(0)$, a convex combination that summarizes each unit’s potential outcome level. Then for the score $g(D,Y,\theta)=HY-\theta$, we have $a(W,\theta)=\bar{Y}$. A simple calculation shows that for $\widehat{\theta}=E_{n}[H_{i}Y_{i}]$ and $\theta_{n}=E_{n}[Y_{i}(1)-Y_{i}(0)]$, we have
$\widehat{\theta}-\theta_{n}=E_{n}[H_{i}a(W_{i})]=\frac{\operatorname{Cov}_{n}(D_{i},\bar{Y}_{i})}{\operatorname{Var}_{n}(D_{i})}.$ (3.2)
Intuitively, the term $E_{n}[H_{i}a(W_{i})]$ from Equation 3.1 isolates the estimator variance due to chance in-sample correlations between the assignments $D_{i}$ and outcome levels $\bar{Y}_{i}$. By contrast, $\phi(W,\theta)=Y(1)-Y(0)-\theta$ does not depend on assignments $D_{i}$, and $\operatorname{Var}(\phi(W,\theta))=\operatorname{Var}(Y(1)-Y(0))$ isolates the estimator variance due to random sampling of units and heterogeneity of individual treatment effects.